# Language-Integrated Recursive Queries

### v1

@@ 2.2.2. Incomplete Results
The query to compute the transitive closure of a directed graph can be expressed in Datalog with two rules: and . This Datalog program is equivalent to the SQL query in Figure 3. However, given a 3-step input graph, i.e. {(0, 1), (1, 2), (2, 3)} this query returns a result with 5 rows (the input edges plus {(0, 2), (1, 3)}) in PostgreSQL (v15) and DuckDB (v1.1). However, when the equivalent query is run in Datalog, the result has 6 rows (the previous results plus (0, 3)). Quickly stepping through the graph shows that 3 is reachable from 0, so why is it not in the RDBMS result?
Unlike Datalog, SQL requires that all WITH RECURSIVE queries be linearly recursive. The way this restriction is often enforced is to limit references to the recursive relation to only once within the recursive subquery. However this is a purely syntactic restriction, so simple aliasing can easily prevent this check from happening, while other databases do not check query linearity at all. Around the time that SQL’99 was preparing to add recursive queries to SQL, there was a belief that “most ‘real life’ recursive queries are indeed linear” (Bancilhon and Ramakrishnan, 1986). While this belief is no longer widely held, this assumption is built into the implementation of many database systems. Within Datalog engines that use the Semi-Naive evaluation algorithm (Ullman, 1989), for the transitive closure example, at each iteration the “new” data is computed with the subquery expressed in the unnamed relational algebra: where represents the data discovered in the previous iteration, and represents all the data discovered so far over the course of iteration. However, if a query is guaranteed to be linear, then the query can be optimized to read only , so the previous query becomes which may cause the algorithm to terminate too early for non-linear queries. Therefore, the mathematical property that leads to incomplete results is non-linearity.
@@
As there are modern databases that do not perform any checks that queries are linear and instead allow the query to return incorrect results, it is of utmost importance that users be prevented from unknowingly sending queries that will silently return invalid results. However recently, MariaDB claimed to add support for non-linear queries (Shalygina and Novikov, 2017), so a user may wish to turn off this constraint depending on their database backend.
2.2.3. Nontermination
@@
WITH RECURSIVE expands the expressive power of SQL to be Turing-complete, so it is possible to represent infinite loops and nontermination. Assuming the SQL specification is followed, that is, all queries are monotonic and linear, what causes nontermination?
Given the query is monotonic over set inclusion, nontermination is the result of infinitely growing intermediate relations. Datalog queries are guaranteed to generate only finite relations, given they are correctly range-restricted over a finite domain (Gottlob et al., 1989). There are a few key differences in the semantics of RDBMS compared to Datalog engines that cause the mathematical properties that provide termination guarantees to be violated.
Range-restriction is defined as the constraint that all variables in the head of a Datalog rule must appear in the body of the rule. Range-restriction applied to SQL queries requires the project clause of each recursive query to contain only constants or references to columns present in relations in the FROM clause. However SQL queries are not limited to a finite domain, as column-level “constructor” operations like addition or string concatenation can arbitrarily introduce new elements. The Datalog literature refers to queries that do not contain operations that can infinitely grow the domain as having the constructor-freedom property (Arntzenius and Krishnaswami, 2016). Finally, the bottom-up evaluation of logic programs assumes set semantics, which provides termination guarantees at the cost of a performance penalty (checking for and removing duplicates). However, RDBMS use bag semantics unless otherwise specified, so a reachability query using bag semantics over data containing cycles will repeatedly re-discover the same paths. Only certain operators like DISTINCT or UNION will use set semantics. For example, given a linearly recursive transitive closure query, and a cyclic input graph, i.e. {(0, 1), (1, 0)}, the query in Figure 3 will infinitely recur. Therefore, the mathematical properties that provide termination guarantees are constructor-freedom and set-semantics.
@@
However, not all queries require set semantics to terminate, although given only the query, it is undecidable whether a query will terminate when using less restrictive duplicate checks (Maher and Ramakrishnan, 1989). If the data does not contain cycles, then the user may prefer to use bag semantics to avoid wasted time on duplicate elimination. Other RDBMS require the use of UNION ALL between the base and recursive case of a recursive query, for example SQLServer and OracleDB, and provide alternative language structures to check for infinite recursion (e.g. CYCLE in OracleDB and PostgreSQL v14+). Users may also wish to turn off the constructor-freedom constraint, as not all constructors lead to infinitely growing domains. Range-restriction, however, will always be enforced. Given the high penalty of an infinitely recurring query, both on the application and the database itself, as other users may also see reduced performance from cross-query interference, it is important to give users the ability to prevent infinitely recurring queries, even at the cost of possible performance overhead or limits to expressibility.
@@ 3. Overview of Embedded Safe Recursion
In Section 2.2, we established the connection between certain problematic database behavior and the mathematical properties of (1) range-restriction, (2) monotonicity, (3) linearity, (4) set-semantics, and (5) constructor-freedom. To avoid generating SQL queries that cause this unwanted behavior, we must design our system to reject queries that do not hold these properties independently.
Figure 4 illustrates how the types representing relations and queries (RestrictedQuery[A, C, D]), and expressions on columns (RestrictedExpr[A, S]) track these properties. Each query is built by applying combinators like map and flatMap on relations. The RestrictedQuery type represents relations that are recursively defined, the A type parameter represents the column-type of the relation, the C type represents the category, e.g. if a query uses bag semantics and can contain duplicates, the D type represents the set of dependency edges between recursive relations. The RestrictedExpr type represents expressions that operate over columns of recursive relations, with A representing the result type and S representing the shape of the expression, e.g. if it contains any non-monotonic operations. Section 4 shows how our system captures these semantic properties of queries and Section 5 uses the host-language type system to translate them into syntactic checks.
The goal of TyQL is not to provide evaluation semantics for recursive queries because database backends vary widely in their interpretation of recursive queries, but there is one consistency across all evaluation engines: the precedence graph of relation dependencies. Therefore, we design our type system to target properties on this graph. Intuitively, each of the five properties can be understood as properties on the program graph of the recursive query. Relations are represented with nodes, dependencies on source relations are represented as directed edges, and recursive queries are subgraphs with cycles. We provide an abstract interpretation that maps terms in to a graph representation in the appendix.
4. Enforcing Mathematical Properties
with
In this section, we present the calculus for recursive queries. We show how the mathematical properties of monotonicity, linearity, set-semantics, range-restriction, and constructor-freedom are encoded into the type system independently so that the restrictions can be composed depending on the targeted database system. Because RDBMS have fragmented support for recursion, yet offer significant performance benefits, it is essential that the type system allow restrictions to be tuned and not overly constrict real-world queries. We first present the DSL type system in Section 4 and then show how to extend it to a 2-level DSL and host language type system in Section 5.
@@ 4.1. Syntax and Base Type System
BOM query
The non-recursive part of extends T-LINQ (Cheney et al., 2013). The syntax of types and terms and the typing judgments are presented in Figure 5. ranges over variables. The column types represent database column types, Named-Tuples represent row types, represents a query that returns a collection of rows of type A, and represents an aggregation that returns a single scalar result of type A. Tuples and function types are used to construct the combinators () that represent database operations. The combinators follow the precedent set by T-LINQ, and we add to represent database-level recursion. We refer to “column-level expressions” as any expression that would go into the filter or project subclause of a SQL query: for map, it would be SELECT a + 1 FROM table vs. for aggregate it would be SELECT sum(a) + 1 FROM TABLE. Both a + 1 and sum(a) + 1 are column-level expressions. Query-level expressions refer to the entire query expression and operations that take relations, e.g. union.
Following the convention set by T-LINQ, we assume a signature that maps each constant c and operator to the typing rule. is useful to abstract over the large set of operations that share the same typing behavior. The operations stored in () can be column-level expressions (, e.g. ); column-level aggregations (, e.g. sum), or relation-level expressions (, e.g. union). Examples of operations stored in are presented in Figure 5.
@@
We build our syntax on the Nested Relational Algebra extended with a “bounded fixpoint” operator (Suciu, 1997). To illustrate an example, in Figure 2 we show a recursive query on the BOM database from Section 2.2 in SQL and . In the TyQL implementation, for-comprehensions desugar to combinators but we do not include for-comprehensions in to keep the syntax minimal. In SQL, a recursive query is composed of a “base case” that contains no references to recursive relations, and a “recursive case” where references to the recursive relation are allowed. Although some databases do not syntactically require a base case, we enforce this in because recursion without a base case leads to undefined behavior. Figure 7 shows the typing rules for the DSL, which closely follows the precedent type system in T-LINQ, with the addition of fix. Judgment states that term has type in the DSL type environment . The fix function defines recursive relations within a single stratum. For , fix takes as arguments a tuple of “base case” queries and a function . The function takes as arguments a tuple of references to the recursive relations being defined and returns a tuple of “recursive case” definitions . Each is composed from the recursive references (and any relations in scope that are defined outside the body of fix) using the combinators in . In the next section, we show how we apply the type system to independently identify and prevent the mathematical properties identified in Section 2.2.
4.2. Range-restricted Recursion
Range-restriction is the constraint that every variable appearing in the head of a rule appears in the body of that rule, to avoid rules that yield infinite relations. Because this constraint does not depend on the semantics of the database system, it is encoded into the basic FIX-D typing rule in Figure 7 by enforcing the recursive references and the recursive definitions all have the same type, and that is a Named-Tuple. As Named-Tuples must have the same key and value types, order, and arity to be considered the same type, this restriction enforces that all variables in the head of a the rule (e.g. columns in project) are present in the body of the rule (e.g. recursive definitions returned by ).
@@ 4.2. Range-restricted Recursion
A function such that holds the range-restricted property if .
4.3. Monotone Recursion
@@
Operations that are non-monotonic under set inclusion are defined as scalar operations. For example, the query SELECT max(a) + 1 FROM TABLE is non-monotonic and produces a single value, while SELECT a + 1 FROM table is monotonic and nonscalar. Both scalar and nonscalar expressions can be constructed from each other: + is nonscalar while sum is scalar but the expressions sum(a + 1) or sum(a) + 1 are both valid and should produce an expression of type scalar, while SELECT x + (SELECT max(y) FROM table1) FROM table2 should produce an expression of type nonscalar (as the subquery does not change that the full expression is monotonic). The type system must be able to distinguish the shape of the entire expression (no matter how nested) so that well-formed terms of type Query are guaranteed to return nonscalar results and are monotonic under set inclusion. The required changes to the type system to restrict recursive queries to monotone operations are shown in Figure 8. We add a type to wrap expressions on columns so that can track the shape () of the expression. We split the OP-D rule into separate rules for , , and . By EXPR-OP-D and AGG-OP-D, any expression that contains an arbitrarily nested sub-expression with type parameter will be of type (via Shape). For example:
The expression on the left contains a nested scalar operation (sum) so the resulting type will be , while the expression on the right contains only nonscalar expressions, so even with the same operator () the result type is The MAP-D and FLATMAP-D rules are refined to only accept functions that contain non-scalar expressions, while AGGREGATE-D and GROUPBY-D only accept functions that contain at least one scalar sub-expression. With these restrictions, expressions of type Query in are guaranteed to contain only monotonic operations.
The monotonicity property is relevant only between the relations in a single stratum (Bancilhon and Ramakrishnan, 1986). To limit the monotonicity restriction to only recursive references, we introduce a RestrictedQuery type and refine FIX-D so that the arguments to must be of type RestrictedQuery. Crucially, RestrictedQuery can only be derived by calling combinators on the arguments passed to the function given to fix as it is not available in the syntax. We update the MAP-D, FLATMAP-D, FILTER-D, and REL-OP-D so if they are passed any arguments of type RestrictedQuery the result type will be RestrictedQuery.
To reduce the number of rules in Figure 8, we define a meta-syntax helper to abstract away the differences between rules for combinators that take RestrictedQuery and Query. Without , we could equivalently define 4 rules for flatMap: if and , the result type will be , however if either or are RestrictedQuery the result will be . As AGGREGATE-D and GROUPBY-D are valid only on Query types, will reject non-monotonic operations on recursive references.
@@
4.4. Linear Recursion

### v2

@@ 3.2. Incorrect Results (B2)
Example. Figure 3 shows an example of a non-linear query, where the recursive relation path is referenced twice in the body of the query. This query is represented by Q2 in Figure 1. The reason why RDBMS may return incorrect results is due to an internal optimization that works only for linear queries and is best illustrated with an example: the query to compute the transitive closure of a directed graph can be expressed in Datalog with two rules: and . This Datalog program is equivalent to the SQL query in Figure 3. Given a 3-step input graph, i.e. {(0, 1), (1, 2), (2, 3)} this query returns a result with 5 rows (the input edges plus {(0, 2), (1, 3)}) in PostgreSQL (v15) and DuckDB (v1.1). However, when the equivalent query is run in Datalog, the result has 6 rows (the previous results plus (0, 3)). Quickly stepping through the graph shows that 3 is reachable from 0, so the result returned from the RDBMS is incorrect.
The reason for this is within Datalog engines that use the Semi-Naive evaluation algorithm, for the transitive closure example, at each iteration the “new” data is computed with the subquery expressed in the unnamed relational algebra: where represents the data discovered in the previous iteration, and represents all the data discovered so far. However, if a query is guaranteed to be linear, then the query can be optimized to read only , so the previous query becomes which may cause the algorithm to terminate too early for non-linear queries. Terminating early may cause the RDBMS to return only partial results, or when nested within an outer query can lead to fully incorrect results.
The SQL specification defines behavior only for WITH RECURSIVE queries that are linearly recursive, leaving the behavior of non-linear queries undefined. Some RDBMS attempt to reject non-linear queries by limiting references to the recursive relation to only once within the recursive subquery. However this is a purely syntactic restriction, so simple aliasing can easily prevent this check from happening, while other databases do not check query linearity at all and allow non-linear queries to execute and silently return incorrect results. Systems that allow mutual recursion and implement it via inlining will show similar behavior and return incorrect results. Therefore, it is the violation of properties P4: linearity or P3: mutual-recursion that leads to behavior B2: incorrect results on systems that allow non-linear or mutual-recursion.
@@
As there are modern databases that do not perform checks for P4 and P3, it is of utmost importance that users be prevented from unknowingly sending queries that will silently return invalid results. However, recently MariaDB claimed to add support for mutual and non-linear queries (Shalygina and Novikov, 2017), so a user may wish to turn off this constraint depending on their RDBMS.
3.3. Nontermination (B3)
@@
WITH RECURSIVE expands the expressive power of SQL to be Turing-complete, so it is possible to express queries that will never terminate. Recursive queries can be computationally expensive, and users may be left unsure if their queries will eventually terminate, given enough time, or never terminate. Assuming the SQL specification is followed, that is, all queries are monotonic and linear, nontermination is a consequence of infinitely growing intermediate relations.
Example. The bottom-up evaluation assumes set semantics (P5), which prevents duplicate tuples from causing the intermediate relations to grow infinitely. Yet RDBMS use bag semantics unless otherwise specified, so a reachability query using bag semantics over data containing cycles will repeatedly re-discover the same paths, leading to infintely growing intermediate relations. Only certain operators like DISTINCT or UNION use set semantics. For example, the query in Figure 3 finds all descendants of a person (‘A’) that are of the same generation (2) over a table Parents(parent, child). Given cyclic input, i.e. {(A, B), (B, A)}, this query will infinitely recur.
Relations can also grow infinitely from non-duplicate tuples if the domain of the query is not finite. SQL queries are not limited to a finite domain, as column-level “constructor” operations like addition or string concatenation can arbitrarily introduce new elements. The query in Figure 3 contains the + operator, which can produce values not in the input domain and can cause nontermination.
Therefore, it is the violation of properties P5: set-semantics, or P6: constructor-freedom that leads to behavior B3: nontermination. On systems that do not reject non-monotonic operations within recursive queries, property P2: monotonicity can also lead to nontermination. Nonterminating queries that violate these properties are represented by Q3 in Figure 1.
@@
Set-semantic recursive queries are guaranteed to generate only finite relations, given they are correctly range-restricted over a finite domain. However, not all queries require set semantics to terminate, although given only the query, it is undecidable whether a query will terminate when using less restrictive duplicate checks (Maher and Ramakrishnan, 1989). If the data does not contain cycles, then the user may prefer to use bag semantics to avoid wasted time on duplicate elimination. Other RDBMS require the use of UNION ALL between the base and recursive case of a recursive query, for example SQLServer and OracleDB, and provide alternative language structures to check for infinite recursion (e.g. CYCLE in OracleDB and PostgreSQL v14+ checks if there is a cycle in the query results). Users may also wish to turn off the constructor-freedom constraint, as not all constructors lead to infinitely growing domains. Given the high penalty of an infinitely recurring query, both on the application and the RDBMS, as other users may also see reduced performance from cross-query interference, it is important to give users the ability to prevent infinitely recurring queries, even at the cost of possible performance overhead or limits to expressibility.
@@ 4. Overview of Embedded Safe Recursion
To avoid generating unsafe SQL queries that cause behaviors B1-B3, we must design our system to reject queries that violate properties P1-P6. Yet, due to the fractured support for recursion in relational databases, as well as the cases where a user may want to relax certain constraints even for a single RDBMS, for our system to be practically and immediately useful it must be possible to pick-and-choose independently and composably which properties should be constrained at any given time. For example, a user of PostgreSQL could choose to constrain mutual-recursion, linearity, and monotonicity but leave the constructor-freedom and set-semantics properties unconstrained. Another user should be able to allow bag semantics for queries where they are certain there are no cycles in the data, and set semantics for when there are cycles. Constraint-independence informs every aspect of the design and evaluation of our approach.
@@
Section 5 describes how the type system captures properties P1-P6 of queries, showing first how each constraint can be implemented independently, then composed into a unified type system. Figure 4 shows an overview of the connection between the types that represent queries and expressions on columns and the targeted properties P1-P6. The A type parameter represents the column-type of the relation, and is used to enforce P1 (range-restriction). For example, a query that returns two columns “age” and “name” will have type (age: Int, name: String). The S type parameter represents the shape of the expression, e.g. if the expression is monotonic or non-monotonic and is used to enforce P2 (monotonicity). The function that constructs RestrictedQuery types, fix, enforces P3 (mutual-recursion). The D type parameter represents the set of dependency edges between recursive relations and is used to enforce P4 (linearity). The C type parameter represents the category, e.g. bag or set semantics, and is used to enforce P5 (set-semantics). The RestrictedExpr type represents expressions that are allowed within recursive queries and is used to enforce P6 (constructor-freedom). Section 6 describes how to translate each restriction into a syntactic check using the host programming language’s type system.
The objective of is not to define a full evaluation semantics for recursive queries, which would require either adopting significant simplifying assumptions or excluding features that are essential for practical applicability. Instead, addresses the more immediate and impactful problem of reasoning about safety in the context of real-world systems, where safety checks on recursive queries are shaped by the needs of users and inconsistent RDBMS implementations. In the appendix, we provide an abstract interpretation that maps terms in to a graph representation that abstracts away differences in RDBMS semantics. This approach enables to apply static safety checks that generalize across engines, without relying on a fixed or idealized semantics.
5. Safe Recursion with
In this section, we present the calculus for recursive queries. We show how the query properties P1-P6 are encoded into the type system independently so that the restrictions can be composed depending on the targeted database system. We first present the DSL type system in Section 5 and then show how to extend it to a 2-level DSL and host language type system in Section 6.
@@ 5.1. Syntax and Base Type System
The base of extends T-LINQ (Cheney et al., 2013). T-LINQ, as well as the Nested Relational Calculus (NRC) (Buneman et al., 1995), established the structural equivalence between SQL-style SELECT-FROM-WHERE and combinator-style flatMap, filter, map, etc. expressions. The recursive query syntax is built on previous work on the NRC with a bounded fixpoint operator (Suciu, 1997).
BOM query
To illustrate an example, Figure 5 shows a recursive query on the BOM database from Section 3 in SQL and . The appendix also includes the queries of Figure 3 in . In TyQL, for-comprehensions desugar to combinators but we do not include for-comprehensions in to keep the syntax minimal. Recursive queries are composed of base cases that contain no references to recursive relations, and recursive cases where references to the recursive relations are allowed. Although some databases do not syntactically require a base case, we enforce this in as recursion without a base case leads to undefined behavior.
@@
The syntax of types and terms and the typing judgments are presented in Figure 6. ranges over variables. Database column types are represented with the column types, multi-column row types with Named-Tuples, represents a relation or query that returns a collection of rows of type A, and represents an aggregation that returns a single scalar result of type A. Tuples and function types are used to construct the combinators () that represent database operations. The combinators follow the precedent set by T-LINQ, and we add fix to represent database recursion. “Column-level expressions” refer to any expression that can go into the filter or project clause of a query: for map, it would be SELECT a + 1 FROM R vs. for aggregate it would be SELECT sum(a) + 1 FROM R. Both a + 1 and sum(a) + 1 are column-level expressions. “Query-level expressions” refer to the full query expression and operations that combine subqueries, e.g. union. The DSL syntax does not include application of functions defined by the user (i.e., ). Function types can be passed to the constructs in but do not get executed until a later stage.
@@
Following the convention set by T-LINQ, we assume a signature that maps each constant c and operator to the typing rule. is useful to abstract over the large set of operations that share the same typing behavior. The operations stored in () can be column-level expressions (, e.g. ); column-level aggregations (, e.g. sum), or relation-level expressions (, e.g. union). In Figure 6 and the typing rules we use the syntax as a stand-in for all operations, even though some operations like or use infix syntax. Examples of operations stored in are presented in Figure 6.
Figure 7 shows the typing rules for the DSL, which closely follows the precedent type system in T-LINQ, with the addition of fix. Judgment states that term has type in the DSL type environment . The fix function defines recursive relations within a single stratum. For , fix takes as arguments a tuple of base case queries and a function . The function takes as arguments a tuple of references to the recursive relations being defined and returns a tuple of recursive case definitions . Each is composed from the recursive references (and any relations in scope that are defined outside the body of fix) using the combinators in . In the next section, we show how we apply the type system to independently identify and prevent properties P1-P6. We start with P1 and include it in the base type system as there are no cases where a user would want to turn off range-restriction.
5.2. P1: Range-restricted Recursion
Range-restriction is the constraint that every variable appearing in the left-hand-side of a Datalog rule appears in the right-hand-side, to avoid rules that yield infinite relations. Because this constraint does not depend on the semantics of the database system, it is encoded into the basic FIX-D typing rule in Figure 7 by enforcing the recursive references and the recursive definitions all have the same type, , and that is a Named-Tuple. As Named-Tuples must have the same key and value types, order, and arity to be considered the same type, this restriction enforces that all variables in the head of a the rule (e.g. columns in project) are present in the body of the rule (e.g. recursive definitions returned by ).
@@ 5.2. P1: Range-restricted Recursion
A function such that holds the range-restricted property P1 if .
5.3. P2: Monotone Recursion
Operations that are non-monotonic under set inclusion are defined as scalar operations. For example, the query SELECT max(a) + 1 FROM R is non-monotonic and produces a single value, while SELECT a + 1 FROM R is monotonic and nonscalar and produces one value per row in R. Both scalar and nonscalar expressions can be constructed from each other: + is nonscalar while sum is scalar but the expressions sum(a + 1) or sum(a) + 1 are both valid and should produce an expression of type scalar, while SELECT x + (SELECT max(y) FROM R1) FROM R2 should produce an expression of type nonscalar (as the subquery does not change that the full expression is monotonic). The type system must be able to distinguish the shape of the entire expression (no matter how nested) so that well-formed terms of type Query are guaranteed to return nonscalar results and hold P2.
@@ 5.3. P2: Monotone Recursion
The required changes to the type system to restrict recursive queries to monotone operations are shown in Figure 8. We add a type to wrap expressions on columns so that can track the shape () of the expression. We split the OP-D rule into separate rules for , , and . By EXPR-OP-D and AGG-OP-D, any expression that contains an arbitrarily nested sub-expression with type parameter will be of type (via Shape). For example, depending on the subexpression, the EXPR-OP-D rule applied to the ”+” operator can produce an expression of type Scalar or NonScalar:
The expression on the left contains a nested scalar operation (sum) so the resulting type will be , while the expression on the right contains only nonscalar expressions, so even with the same operator () the result type is . The MAP-D and FLATMAP-D rules are refined to only accept functions that return non-scalar expressions, while AGGREGATE-D and GROUPBY-D only accept functions that return at least one scalar sub-expression. With these restrictions, expressions of type Query are guaranteed to contain only monotonic operations, while only expressions of type Aggregation may contain non-monotonic operations.
As discussed in Section 2, the monotonicity property only needs to restrict aggregation between the relations in a single stratum, i.e., within recursive queries to recursive references. To limit the monotonicity restriction to only recursive references, we introduce a RestrictedQuery type and refine FIX-D so that the arguments and return type of must be of type RestrictedQuery. Crucially, RestrictedQuery can only be derived by calling combinators on the arguments passed to the function given to fix as it is not available in the syntax and therefore can only be in scope within the body of fix. We update the MAP-D, FLATMAP-D, FILTER-D, and REL-OP-D so if they are passed any arguments of type RestrictedQuery the result type will be RestrictedQuery.
To reduce the number of rules in Figure 8, we define a meta-syntax helper to abstract away the differences between rules for combinators that take RestrictedQuery and Query. Without , we could equivalently define 4 rules for flatMap: if and , the result type will be , however if either or the return type of are RestrictedQuery the result will be . As AGGREGATE-D and GROUPBY-D are valid only on Query types, aggregations cannot be applied to recursive references and will reject non-monotonic operations on recursive relations within recursive queries.
@@
5.4. P3: Mutually-recursive Recursion
restricts mutually-recursive queries by simply limiting the number of arguments to the fix function to a single argument. Single-direction dependencies between recursive relations can be defined using multiple calls to fix.
Definition 3.
A function such that holds the mutually-recursive property P3 if .
5.5. P4: Linear Recursion